OpenStudy (anonymous):

x^2+8x+c... what is the vertex?

OpenStudy (anonymous):

x = -4 and y= (-4)^2 + 8(-4) +c = 16-32 +c = -16+c

OpenStudy (anonymous):

The vertex is (-4,-16+c)

OpenStudy (anonymous):

how'd u get -4 tho